# CSC-MPPI: A Novel Constrained MPPI Framework with DBSCAN for Reliable Obstacle Avoidance*

### v1

@@ V CONCLUSIONS
In this paper, we introduced a novel Constrained Sampling Cluster MPPI (CSC-MPPI) framework that integrates a primal-dual gradient-based constraint enforcement method with DBSCAN clustering to overcome the inherent limitations of standard MPPI in satisfying hard constraints. By iteratively adjusting sampled control inputs into the feasible region and clustering them based on spatial and cost similarities, the proposed approach effectively selects the optimal control action while strictly adhering to both state and input constraints. Simulation and real-world experimental results demonstrate that CSC-MPPI achieves a zero collision rate and generates shorter, more feasible trajectories compared to conventional MPPI methods, thereby validating its effectiveness and robustness in complex obstacle avoidance tasks. Future work will focus on enhancing computational efficiency and extending the framework to more challenging environments and diverse robotic platforms including a mobile manipulator and humanoid.

### v2

@@ V CONCLUSIONS
In this paper, we introduced a novel CSC-MPPI framework that integrates a primal-dual gradient-based constraint enforcement method with DBSCAN clustering to overcome the inherent limitations of standard MPPI in satisfying hard constraints. By iteratively adjusting sampled control inputs into the feasible region and clustering them based on spatial and cost similarities, the proposed approach effectively selects the optimal control action while strictly adhering to both state and input constraints. Simulation and real-world experimental results demonstrate that CSC-MPPI achieves a zero collision rate and generates shorter, more feasible trajectories compared to conventional MPPI methods, thereby validating its effectiveness and robustness in complex obstacle avoidance tasks. Future work will focus on enhancing computational efficiency and extending the framework to more challenging environments and diverse robotic platforms including a mobile manipulator and humanoid.